OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = 2x2 – 8x + 6? x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 3) x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 6) x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 3) x–intercepts: (3, 0) and (1, 0) y–intercept: (0, 6)

OpenStudy (anonymous):

0 = 2x^2 - 8x + 6 0 = x^2 - 4x + 3 (x-3) (x -1) = 0 x = 3, 1

OpenStudy (anonymous):

To find the x-intercept, just set y=0 and solve for X. Similarly, to find the y-intercept, set x=0 and solve for Y.

OpenStudy (anonymous):

x–intercepts: (3, 0) and (1, 0) y–intercept: (0, 6)